# The Schwurbelarchiv: a German Language Telegram dataset for the Study of Conspiracy Theories

### v2

@@ Abstract
Sociality borne by language, as is the predominant digital trace on text-based social media platforms, harbours the raw material for exploring multiple social phenomena. Distinctively, the messaging service Telegram provides functionalities that allow for socially interactive as well as one-to-many communication. Our Telegram dataset contains over 6,000 groups and channels, 40 million text messages, and over 3 million transcribed audio files, originating from a data-hoarding initiative named the “Schwurbelarchiv” (from German schwurbeln: speaking nonsense). This dataset publication details the structure, scope, and methodological specifics of the Schwurbelarchiv, emphasising its relevance for further research on the German-language conspiracy theory discourse. We validate its predominantly German origin by linguistic and temporal markers and situate it within the context of similar datasets. We describe process and extent of the transcription of multimedia files. Thanks to this effort the dataset uniquely supports multimodal analysis of online social dynamics and content dissemination. Researchers can employ this resource to explore societal dynamics in misinformation, political extremism, opinion adaptation, and social network structures on Telegram. The Schwurbelarchiv thus offers unprecedented opportunities for investigations into digital communication and its societal implications.
Keywords telegram dataset, conspiracy theories, social media, misinformation
@@ 1 Introduction
The proliferation of digital trace data from social media platforms has enriched the study of societal phenomena, enabling researchers to analyze communication patterns and the dissemination of information at large scales Freelon (2014). Platforms like Telegram, with its blend of interactive groups and broadcast channels, serve as fertile grounds for examining issues ranging from political discourse to the spread of conspiracy theories. The insights gained from such analyses reveal their influence on societal narratives.
@@
The Schwurbelarchiv dataset contributes to a growing number of published Telegram datasets for research, e.g. La Morgia et al. (2023); Baumgartner et al. (2020), and – with the transcription of audio files – offers a unique starting point to widen the scope of research into conspiracy theories. This publication aims to provide a resource for researchers to enable large-scale investigations of the societal and linguistic dynamics that are connected with conspiracy theories, misinformation, political dissatisfaction, and extremism online, which has become a growing research interest in recent years Urman and Katz (2022); Curley et al. (2022); Bovet and Grindrod (2022); Hoseini et al. (2023); Willaert et al. (2022); Zehring and Domahidi (2023). The following paper first gives insights into scope and creation process of the dataset with an emphasis on transcribed audio data, and an estimation of how much of the German-language conspiracy-related discourse on Telegram is covered in our dataset.
1.0.1 Telegram.
@@
Telegram is an instant messaging platform where users can exchange messages by creating an account with their phone number. Users can communicate in direct chats (one-to-one communication), in groups of up to 200,000 users (many-to-many), or participate in broadcast channels (one-to-many). These broadcast channels allow an unlimited amount of Telegram users to follow a single Telegram user’s (the administrator’s) content, and therefore broadcast channels are well designed for information dissemination at a large scale. While one-to-one communication is private and generally not accessible to researchers, groups and channels are public and accessible if their URL is known. We refer to groups and channels together as “chats” in the remainder of this work.
@@ 2 Related Works
A number of studies on Telegram datasets or research primarily using Telegram data have been conducted in recent years, many of which focus on research on the dissemination of misinformation and conspiracy theories in online environments. Here, we introduce the three Telegram datasets most closely related to our work in terms of temporal and linguistic scope of the data contained. We analyse similarities and differences across datasets, aiming to provide an assessment of the coverage of the Schwurbelarchiv and situating the dataset within the relevant research context.
2.0.1 Public Telegram Datasets.
@@
The study by  Zehring and Domahidi (2023) investigates the connections between the Querdenken movement, a prominent anti-COVID protest network in German-speaking countries often linked to conspiracy theories, to far-right and alternative media, focusing on content-sharing patterns and the influence of QAnon within the network. Using a dataset of 6,294,955 text messages from 578 German public Telegram channels, the research explores the movement’s key topics over time and across its sub-communities. The dataset was collected by first identifying seed channels by searching for actors connected to the Querdenken and far-right movements, to the esoteric/spirituality scene as well as nominees for the Golden Tin Foil Hat prize  [^1] https://www.dergoldenealuhut.de in Germany, Austria, and Switzerland. This approach resulted in 238 initial channels, from which all messages were scraped. Next, frequently forwarded channels (mentioned 50 times or more) were identified and included in the dataset. This procedure – also called “snowball sampling” – was performed twice and led to a total of 578 public Telegram channels from which all messages available for the time between October 28, 2015 and January 3, 2022 were collected. By performing topic modelling on the dataset, Zehring and Domahidi (2023) found four overarching topics present in the Telegram channels: promotion, QAnon, right-wing populism, and COVID-19 conspiracy theories.
@@
A larger (at the time of writing unpublished) dataset on similar topics was assembled by Mohr (2023). Since snowball sampling on its own quickly yields an unmanageable amount of chats, they implemented a guidance mechanism to prioritize chats likely to yield relevant data by predicting their relevance based on a keyword list which was translated into 44 languages. The keywords and therefore selection of chats was targeted toward content containing misinformation related to COVID-19. The dataset which is to our knowledge one of the largest Telegram datasets available, includes 433,000 chats, with full data from 128,000 chats comprising 2.1 billion messages, primarily in Russian, English, and German. Messages, spanning text, metadata, and links (but excluding videos, images, and voice messages), predominantly reflect activity from European time zones and trends linked to events like the European COVID-19 vaccination campaign and the start of the Russian invasion of Ukraine in the time period of January 2020 to January 2023. All predominantly German language chats (23,000) from this dataset were made available to the authors of this study for comparison to the Schwurbelarchiv.
@@
The dataset collected by Baumgartner et al. (2020), the largest publicly available collection of Telegram data to date, includes 27,000 chats and 317 million messages. The dataset is available for the years 2015-2019 and was regularly updated during that time, with source code available to facilitate further data collection, enabling researchers to extract information from specific chats. Starting with a seed list of 250 English-language chats – 124 focused on right-wing extremism and 137 on cryptocurrency – the number of chats included was also expanded using a snowball sampling approach. All of the studies described here used the Telegram API to build customized clients and collect data.
2.0.2 Message Deletion.
The deletion of messages by users on Telegram has emerged as a significant factor influencing the completeness and reliability of Telegram datasets. The study by Buehling (2024) examines the phenomenon of message deletion in Telegram chats, focusing on its implications for computational social science. Using a repeatedly scraped dataset of messages from 25 groups and 25 channels, the study highlights how the nature of digital trace data can lead to systematic biases in analyses such as topic modelling, dictionary-based computational content analysis, and network studies. The study finds that for channels, on average, only 88% of messages remain accessible five days after posting, dropping to 83% after seven months. In groups, the decline is even steeper, with only 64% of messages remaining after five days and 52% after seven months, emphasizing the need for timely data collection. Furthermore, Buehling (2024) finds that the types of messages deleted vary substantially between channels and groups. In groups, text messages are disproportionately deleted, while in channels, forwarded content and videos are more frequently removed. These discrepancies introduce biases not only in computational content analyses but also in the mapping of information flows and actor networks and can skew conclusions drawn about the nature and dynamics of online communication.
@@ 3 Methodology
The entire raw data contained in the Schwurbelarchiv is available on the Internet Archive [^2] https://archive.org/details/schwurbel-archiv [^3] https://web.archive.org/web/20241126233137/https://schwurbelarchiv.wordpress.com/. The person initially responsible for collecting the data remains anonymous, but responds to requests via email and provides information about the data selection process. They describe their motivation for collecting the dataset as the pure joy of data-hoarding, contribution to the Internet Archive, and enablement of research and journalism using the collected data as a source.
@@
Similar to the other Telegram datasets described here, groups and channels in the Schwurbelarchiv were selected via a snowball sampling approach. Starting with a seed set of Telegram chats focused on renowned German speaking conspiracy theorists, the process involved the automated isolation of invitation links to other Telegram chats and joining the chats to collect the messages posted there. Relevance of chats identified this way, e.g. a focus on conspiracy-related content, was determined by a human-in-the-loop to limit the number of included chats. After identification of a relevant chat, its content was initially downloaded in bulk, after which a continual streaming process for content posted in the chat was implemented. In contrast to the data collection approach of the other studies described here, the data was gathered using a Windows Virtual Machine, with four separate Telegram Desktop sessions, each operating under its own account. Chats were then exported using the Remote Desktop Protocol (RDP). A Python script handled the zipping of files to optimize storage, extracted URLs, and uploaded the data. Local files were removed after the upload. RDP enables storing not only the HTML text files, but all video files, voice messages, and images posted in the chats. Video files larger than 8 MB have not been downloaded until mid of 2021, whereafter the limit for downloaded videos was increased to 50 MB. To our knowledge, the Schwurbelarchiv is the only publicly available Telegram dataset that contains multimedia content. This content constitutes a large part of the discourse taking place on social media platforms that is typically ignored due to the difficulty of handling diverse data formats and large data volumes. The fact that our dataset retains this content greatly expands the types of research questions and analysis approaches that it enables.
3.0.2 Data Cleaning and Preprocessing
@@
Since the raw collected data contains multimedia files, it requires a storage space of nearly 24 TB in a compressed state. The sheer size of the data as well as the native Internet Archive file format .warc complicate processing and analysis. Therefore, next to clarifying ethical and licensing questions associated with the use of the Schwurbelarchiv for research (see Section “Ethics and Data Protection” below), we set out to preprocess the dataset to make it more easily accessible for a wider research community. This preprocessing involved data cleaning, transcription of audio content, and the validation of the completeness of the dataset. We describe these steps in more detail in the remainder of this section.
@@
After downloading the raw data from the Internet Archive, we used Python scripts to extract all relevant information, namely author, posting date, and message body. In addition, we distinguish between original messages and forwarded messages – and their respective authors and posting dates – which enables the reconstruction of message forwarding networks within the dataset. Furthermore, we isolate links to websites shared in the message text body. A full listing of all data columns contained in the cleaned dataset is provided in the appendix 5. While the raw dataset as it is published on the Internet Archive contains uncensored usernames, we pseudonymise usernames using a sha-256 hash-function. We extract usernames from the HTML files containing the content of the chats and do not have access to user IDs. As a result, we cannot distinguish between users that have chosen the same username. This is a substantial limitation of our dataset.
@@
During the initial data collection, there was no quality control of the included chats with respect to research interests such as the investigation of conversation patterns. However, the raw dataset contains a substantial number of “dumps” – chats that were started anew by dumping all the content of a discontinued chat into them. Such chats add redundant information to the dataset, might artificially skew the relation of content posted per author, and distort analyses that leverage the time stamps of messages to reconstruct temporal patterns. To remove such dumps, we therefore excluded chats that either showed activity for a period shorter than 10 days or where more than 90% of the content included was posted in a single day. In total, we excluded 527 such dump chats with 163,056 messages this way. While we do not further use the content from these chats, their sheer existence provides an estimator of how many chats in the domain of conspiracy discussions on Telegram are discontinued. After cleaning the dataset, it contains a total of 56,230,491 messages posted in 6067 chats, covering the time span from September 23, 2015 to August 5, 2022 (see also Tab. 1).
3.0.3 Networks
In addition to the analysis of the content of messages within groups, the Schwurbelarchiv allows for the analysis of the network structure of group affiliation of authors. Viewed through the lens of networks, our dataset can be conceptualised as a network in which groups are nodes and a weighted edge between two nodes indicates how many authors did at least post one message in both groups. In a preliminary analysis, by calculating average degree and edge weight, we find that authors have potential outreach to a mean of 2522.4 other authors (median = 637.0), and share 3.49 common groups with other authors (median = 3.32). Likewise, the data set can be conceptualized as a network in which authors are nodes and a weighted edge between two authors indicates in how many groups both authors participated in the discourse. Here it shows, that a groups shares common at least one common author with a mean average of 1289.0 other groups (median = 1457.0), and groups share a mean of 8.16 authors (median = 2.19). Both networks make it possible to examine topological features of posting activity in the dataset. Based on the messages contained in the Schwurbelarchiv we compiled these networks and make them available for further analysis.
3.0.4 Ethics and Data Protection

### v3

# The Schwurbelarchiv: A German-Language Dataset Derived from Multimodal Telegram Content for the Study of Conspiracy Theories

@@ Abstract
Sociality borne by language, as is the predominant digital trace on text-based social media platforms, harbours the raw material for exploring a multitude of social phenomena. Distinctively, the messaging service Telegram provides functionalities that allow for socially interactive as well as one-to-many communication. Our Telegram dataset contains over 5,800 groups and channels and 63 million messages, originating from a data-hoarding initiative named the “Schwurbelarchiv” (from German schwurbeln: speaking nonsense). Uniquely, it includes the transcriptions of over 3 million audio and video files. While the raw data was previously archived on the Internet Archive by an anonymous data hoarder, it was stored in a format that is difficult to process and largely inaccessible for systematic research. Our contribution consists of parsing, cleaning, and validating this raw archive, pseudonymising user data, and transcribing roughly 126,000 hours of audio and video content, thereby transforming this data hoard into a structured, research-ready dataset. This dataset publication details the structure, scope, and methodological specifics of the Schwurbelarchiv, emphasising its relevance for further research on the German-language conspiracy-theory-related discourse. We validate its predominantly German origin by linguistic and temporal markers and situate it within the context of similar datasets. We describe process and extent of the transcription of multimedia files. Thanks to this effort the dataset uniquely supports analysis of text from originally multimodal sources like voice messages and videos to investigate online social dynamics and content dissemination. Researchers can employ this resource to explore societal dynamics related to misinformation, political extremism, opinion adaptation, and social network structures.
keywords:
telegram dataset, conspiracy theories, social media, misinformation
@@ 1 Introduction
The proliferation of digital trace data from social media platforms has enriched the study of societal phenomena, enabling researchers to analyse communication patterns and the dissemination of information at large scales [1]. Platforms like Telegram, with its blend of interactive groups and broadcast channels, serve as fertile ground for examining issues ranging from political discourse to the spread of conspiracy theories. The insights gained from such analyses reveal their influence on societal narratives.
@@
The Schwurbelarchiv [^1] The term schwurbeln, while denoting pseudoscientific and irrational content more generally, gained widespread usage in German-speaking countries from 2021 onward in the context of the COVID-19 pandemic. dataset introduced here contributes to a growing number of published Telegram datasets for research, e.g. La Morgia et al. [2], Baumgartner et al. [3], and – with the transcription of audio files – offers a unique starting point to widen the scope of research into conspiracy theories by including multimedia content. This publication aims to provide a resource for researchers to enable large-scale investigations of the societal and linguistic dynamics that are connected with conspiracy theories, misinformation, political dissatisfaction, and extremism online, which has become a growing research interest in recent years [4, 5, 6, 7, 8, 9].
For empirical research on conspiracy theories in German-language in particular, digital trace data from Social Media Platforms or Messenger Services have substantially contributed to the research landscape but still face limitations that we aim to address with the data set described here.  Pustet et al. [10] study conspiracy content in approximately 4,000 German Telegram messages and identify the scarcity of large-scale German data as a limitation for the robustness of the research and its applicability to the broader discourse. They also note that existing datasets are predominantly in English and often keyword-filtered instead of snowball-sampled, which creates systematic biases towards a few particular conspiracy theories and narratives. Next to limitations of scale and sampling, the exclusive reliance on textual data leaves a substantial portion of Telegram communication unstudied. In our dataset, media files such as images, videos, and voice messages make up 24.4% of all messages.  Hartwig et al. [11] explicitly identify the neglect of voice messages in misinformation research as a critical gap, pointing to the absence of accessible voice message databases despite the growing role of audio content in the dissemination of conspiracy theories on platforms like Telegram and WhatsApp. Similarly, Steffen [12] takes a first step towards multimodal analysis of German conspiracy channels by applying topic modelling to text-image pairs, yet notes that audio content remains entirely outside the scope of current approaches, despite its prevalence on Telegram. The Schwurbelarchiv addresses these limitations by providing a large-scale, snowball-sampled, German-language dataset of over 63 million messages from more than 5,800 Telegram chats, combined with the transcription of roughly 126,000 hours of audio and video content, which allows text analyses of multimodal sources that move beyond the text-only, small-scale, and English-centric resources currently available.
We emphasise that the raw data contained in the Schwurbelarchiv was not collected by the authors of this paper but was assembled by an anonymous individual and deposited on the Internet Archive. In its original form, the archive comprises nearly 24 TB of compressed multimedia files in formats that are not readily accessible for computational analysis. The contribution of the our work is therefore not the data collection itself, but the substantial effort of making this data usable for research: we parse and clean the raw archive, pseudonymise user information, assess the dataset’s coverage, and transcribe over three million audio and video files. The result is a structured, text-based dataset that can be directly used for computational social science research.
1.1 Telegram.
@@
Telegram is an instant messaging platform where users can exchange messages by creating an account with their phone number. Users can communicate in direct chats (one-to-one communication), in groups of up to 200,000 users (many-to-many), or participate in broadcast channels (one-to-many). These broadcast channels allow an unlimited amount of Telegram users to follow a single Telegram user’s (the administrator’s) content, and therefore broadcast channels are well designed for information dissemination at a large scale. While one-to-one communication is private and generally not accessible to researchers, groups and channels are public and accessible if their URL is known. We refer to groups and channels together as “chats” in the remainder of this work. Notably, Telegram does not offer a global search function for discovering chats by topic, which necessitates snowball sampling approaches as employed by the individual that curated the dataset described here.
@@ 2 Related Works
A number of studies on Telegram datasets or research primarily using Telegram data have been conducted in recent years, many of which focus on the dissemination of misinformation and conspiracy theories in online environments. Here, we introduce the three Telegram datasets most closely related to our work in terms of temporal and linguistic scope of the data contained. We analyse similarities and differences across datasets, aiming to provide an assessment of the coverage of the Schwurbelarchiv and situating the dataset within the relevant research context.
2.1 Public Telegram Datasets.
@@
The study by  Zehring and Domahidi [9] investigates the connections between the Querdenken movement, a prominent anti-COVID protest network in German-speaking countries often linked to conspiracy theories, to far-right and alternative media, focusing on content-sharing patterns and the influence of QAnon within the network. Using a dataset of 6,294,955 text messages from 578 German public Telegram channels, the research explores the movement’s key topics over time and across its sub-communities. The dataset was collected by first identifying seed channels by searching for actors connected to the Querdenken and far-right movements, to the esoteric/spirituality scene as well as nominees for the Golden Tin Foil Hat prize  [^2] https://www.dergoldenealuhut.de in Germany, Austria, and Switzerland. This approach resulted in 238 initial channels, from which all messages were scraped. Next, frequently forwarded channels (mentioned 50 times or more) were identified and included in the dataset. This procedure – also called “snowball sampling” – was performed twice and led to a total of 578 public Telegram channels from which all messages available for the time between October 28, 2015 and January 3, 2022 were collected. By performing topic modelling on the dataset, Zehring and Domahidi [9] found four overarching topics present in the Telegram channels: promotion, QAnon, right-wing populism, and COVID-19 conspiracy theories.
@@
A larger, at the time of writing unpublished, dataset on similar topics was assembled by Mohr [13]. Since snowball sampling on its own quickly yields an unmanageable amount of chats, they implemented an automated guidance mechanism that prioritises chats likely to contain relevant data based on a keyword list translated into 44 languages. The keywords targeted content containing misinformation related to COVID-19. The dataset, which is to our knowledge of the largest Telegram datasets available, includes 433,000 chats, comprising 2.1 billion messages, primarily in Russian, English, and German. The dataset contains text, metadata, and links, but excludes videos, images, and voice messages, which is a limitation that is central to the distinction with the Schwurbelarchiv. The messages predominantly reflect activity from European time zones in the period from January 2020 to January 2023, with trends linked to events such as the European COVID-19 vaccination campaign and the start of the Russian invasion of Ukraine. All predominantly German-language chats (23,000) from this dataset were made available to the authors of this study for comparison to the Schwurbelarchiv.
@@
The dataset collected by Baumgartner et al. [3], the largest publicly available collection of Telegram data to date, includes 27,000 chats and 317 million messages. The dataset is available for the years 2015-2019 and was regularly updated during that time, with source code available to facilitate further data collection, enabling researchers to extract information from specific chats. Starting with a seed list of 250 English-language chats – 124 focused on right-wing extremism and 137 on cryptocurrency – the number of chats included was also expanded using a snowball sampling approach. As for the dataset by Mohr [13], this data set includes textual data only. Table 1 provides a summary comparison of these datasets.
Schwurbelarchiv Zehring & Domahidi [9] Mohr [13] Pushshift [3] TGDataset [2]
Chats 5,807 578 22,968 27,801 120,979
Messages 63M 6.3M 2.1B∗∗ 317M 498M
Time span 2015–2022 2015–2022 2020–2023 2015–2019 2015–2022
Language German German Multilingual Multilingual Multilingual
Topic focus Conspiracy Conspiracy / far-right COVID misinfo General General
Transcriptions Yes (126,839 h) No No No No
Publicly available Upon request No No Yes Yes
Table 1: Comparison of Telegram datasets. German-language subset; the full dataset contains 128,000 chats. ∗∗Total across all languages. Channels only; no groups.
2.2 Research on German-Language Conspiracy Content on Telegram.
Beyond the collection of datasets, a number of studies have used German-language Telegram data to investigate conspiracy-related content using digital-trace data and computational methods.  Weigand et al. [14] performed a long-term content analysis of conspiracy narratives in the German Querdenken protest movement, fine-tuning a DistilBERT model on self-annotated data from 34 chat groups comprising 822,000 text messages.  Pustet et al. [10] compared supervised fine-tuning of BERT-based models with prompt-based approaches using GPT-4 and Llama 2 for detecting conspiracy theories in German Telegram messages, building on an annotated corpus of approximately 4,000 messages. More recently,  Steffen [12] applied topic modelling on text-image pairs from 571 German conspiracy channels but still ignores voice messages and videos. These studies share a common set of limitations that our dataset helps to overcome: They operate on comparatively small corpora and are limited to text or at most text and images. Audio and video content constitute a considerable share of Telegram messages, but has not been incorporated into any of these analyses. The Schwurbelarchiv, with its scale and the transcription of audio files, is designed to overcome several of these limitations. For instance, the dataset enables investigating how conspiracy narratives evolve across thousands of groups over time, examining the role of voice messages and videos in disseminating conspiracy content compared to text, or tracing information flows through forwarding networks to understand how specific narratives spread between groups.
@@ 3 Methodology
The entire raw data contained in the Schwurbelarchiv is available on the Internet Archive [^3] https://archive.org/details/schwurbel-archiv [^4] https://web.archive.org/web/20241126233137/https://schwurbelarchiv.wordpress.com/. The person initially responsible for collecting the data remains anonymous, but responded to requests via email by the authors and provided information about the data selection process. They describe their motivation for collecting the dataset as the pure joy of data-hoarding, contribution to the Internet Archive, and enablement of research and journalism using the collected data as a source.
Similar to the other Telegram datasets described here, groups and channels in the Schwurbelarchiv were selected via a snowball sampling approach. Starting with a seed set of Telegram chats focused on renowned German-speaking conspiracy theorists, other chats were identified via invitation links and forwarded messages originating from the chats in the sample. Newly discovered chats were added to the sample, and all messages posted in them up to the point of their discovery were scraped via bulk download, after which a continual streaming process captured new content in near real-time. Relevance of newly discovered chats was determined by manual curation (e.g., human in the loop) that excluded chats deemed unrelated to conspiracy theories in order to prune irrelevant sampling paths early. Data collection was terminated when the anonymous collector lacked the resources to continue. The resulting dataset covers the time span from September 23, 2015 to August 5, 2022.
Snowball sampling promises to capture a more representative portion of the discourse but still has a number of limitations: manual curation of discovered chats ensures relevance but remains subjective, relevant chats might not be included because sampling terminated before a link to them was found, and relevant chats might be missed because they are entirely disconnected from the forwarding network in which sampling was initiated. In particular, chats from which few messages are forwarded are systematically less likely to be discovered by this approach. We note, however, that snowball sampling is the only available methodology for data collection from Telegram beyond manual search and selection, as the platform lacks a searchable index of public chats.
In contrast to the data collection approach of the other studies described here which relied on the Telegram API, the data in the Schwurbelarchiv was gathered using a Windows Virtual Machine, with four separate Telegram Desktop sessions, each operating under its own account. Chats were then exported using the Remote Desktop Protocol. Once a chat has been discovered by the snowball sampling process, all it’s content up to that point in time was bulk-downloaded. Notably, this approach also enables streaming of the content after a chat has been discovered. This is important because the deletion of messages by users on Telegram has emerged as a significant factor influencing the completeness and reliability of Telegram datasets. Buehling [15] highlight how this can lead to systematic biases in analyses such as topic modelling, dictionary-based computational content analysis, and network studies. The study finds that for channels, on average, only 88% of messages remain accessible five days after posting, dropping to 83% after seven months. In groups, the decline is even steeper, with only 64% of messages remaining after five days and 52% after seven months, emphasizing the need for timely data collection.
@@
The Rempote Desktop Protocol enables storing not only the HTML text files, but all video files, voice messages, and images posted in the chats. Video files larger than 8 MB have not been downloaded until mid of 2021, whereafter the limit for downloaded videos was increased to 50 MB. To our knowledge, the Schwurbelarchiv is the only publicly available Telegram dataset that contains multimedia content. This content constitutes a large part of the discourse taking place on social media platforms that is typically ignored due to the difficulty of handling diverse data formats and large data volumes. The fact that our dataset retains this content and makes the multimodal files accessible through transcribing them into text-data greatly expands the types of research questions and analysis approaches that it enables.
All data columns describing message content and metadata (see Table 5) were parsed from the HTML files exported by the original data collector. The columns message_lang, audio_lang, and video_lang were added by us using the Lingua language detection package  [^5] https://github.com/pemistahl/lingua-py, is_conspiracy was computed using the fine-tuned BERT model by Pustet et al. [10] as described in Section “Conspiracy Content in Telegram Chats” below, and all transcription-related columns (audio_transcription, video_transcription, and associated duration and file size fields) were generated by us using the Whisper model described in the following section.
3.2 Data Cleaning and Preprocessing
@@
Since the raw collected data contains multimedia files, it requires a storage space of nearly 24 TB in a compressed state. The sheer size of the data as well as the native Internet Archive file format .warc complicate processing and analysis. Therefore, next to clarifying ethical, data protection-related, and licensing questions associated with the use of the Schwurbelarchiv for research (see Section “Ethics and Data Protection” below), we set out to preprocess the dataset to make it more easily accessible for a wider research community. This preprocessing involved data cleaning, transcription of audio content, and an assessment of the completeness of the dataset. We describe these steps in more detail in the remainder of this section. The fully documented pipeline is available on GitHub [^6] https://github.com/cs2-lab/schwurbelarchiv.
@@
After downloading the raw data from the Internet Archive, we used Python scripts to extract all relevant information, namely author, posting date, and message body. In addition, we distinguish between original messages and forwarded messages – and their respective authors and posting dates – which enables the reconstruction of message forwarding networks within the dataset. Furthermore, we isolate links to websites shared in the message text body. A full listing of all data columns contained in the cleaned dataset is provided in the Appendix 5. While the raw dataset as it is published on the Internet Archive contains uncensored usernames, we pseudonymize usernames using a SHA-256 hash-function. We extract usernames from the HTML files containing the content of the chats but do not have access to user IDs assigned by Telegram. As a result, we cannot distinguish between users that have chosen the same username. This is a substantial limitation of our dataset. We further note that Telegram allows the use of bots as automated accounts, which may also be present in the dataset. We did not attempt to identify or flag bot accounts, which constitutes another limitation of the dataset.
In addition to textual data, we transcribed the audio of both audio and video files. We used ffmpeg to first transcode the various audio codecs into WAV, and openai/whisper-large-v3-turbo [^7] https://huggingface.co/openai/whisper-large-v3-turbo to transcribe the audio content. Various optimizations were used in order to make the transcription task more computationally efficient, including using the CTranslate2 faster-whisper implementation [^8] https://github.com/SYSTRAN/faster-whisper with Voice Activity Detection and batching. These resulted in a transcription real time factor of $\sim 250$X per Nvidia L40S GPU. Overall, 126,839 hours of content (14.9 years) were transcribed.
@@
During the initial data collection, there was no quality control of the included chats with respect to research interests such as the investigation of temporal conversation patterns. However, the raw dataset contains a substantial number of “dumps” – chats that were started anew by dumping all the content of a discontinued chat into them. Such chats add redundant information to the dataset, might artificially skew the relation of content posted per author, and distort analyses that leverage the time stamps of messages to reconstruct temporal patterns. To remove such dumps, we therefore excluded chats that either showed activity for a period shorter than 10 days or where more than 90% of the content included was posted in a single day. In total, we excluded 1,050 such dump chats (13.7% of total) with 317,532 messages (0.5% of total). While we do not further use the content from these chats, their sheer existence provides an estimation of how many chats in the domain of conspiracy discussions on Telegram are discontinued. After cleaning the dataset, it contains a total of 63,908,365 messages posted in 5807 unique chats, covering the time span from 23/09/2015 to 05/08/2022 (see also Tab. 2).
3.3 Ethics and Data Protection